# Discrete Neural Flow Samplers with Locally Equivariant Transformer

### v1

@@ 1 Introduction
Building on these advances, the goal of our paper is to develop a sampling method for discrete distributions that is both efficient and scalable. To this end, we introduce Discrete Neural Flow Samplers (DNFS), a novel framework that learns the rate matrix of a CTMC whose dynamics satisfy the Kolmogorov forward equation (Oksendal, 2013). In contrast to discrete flow models (Campbell et al., 2024; Gat et al., 2024), which benefit from access to training data to fit the generative process, DNFS operates in settings where no data samples are available. This data-free setting makes direct optimisation of the Kolmogorov objective particularly challenging and necessitates new methodological advances to ensure stable and effective training. Specifically, the first difficulty lies in the dependence of the objective on the intractable partition function. We mitigate this by using control variates (Geffner & Domke, 2018) to reduce the variance of its Monte Carlo estimate, which enables efficient optimisation via coordinate descent. More critically, standard neural network parameterisations of the rate matrix render the objective computationally prohibitive. To make learning tractable, a locally equivariant Transformer architecture is introduced to enhance computational efficiency significantly while retaining strong model expressiveness. Empirically, DNFS proves to be an effective sampler for discrete unnormalised distributions. We further demonstrate its versatility in diverse applications, including training discrete energy-based models and solving combinatorial optimisation problems.
@@ 2 Preliminaries
In the next section, we describe how to leverage Equation 4 to learn a model-based rate matrix for sampling from a given target distribution $\pi$, followed by the discussion of applications to discrete energy-based modelling and combinatorial optimisation.
@@ 3 Discrete Neural Flow Samplers
The rate matrix $R_{t}$ that transports the initial distribution to the target is generally not unique. However, we can select a particular path by adopting an annealing interpolation between the prior $\eta$ and the target $\rho$, defined as $p_{t}\propto\rho^{t}\eta^{1-t}\triangleq\tilde{p}_{t}$ (Gelman & Meng, 1998; Neal, 2001). This annealing path coincides with the target distribution $\pi\propto\rho$ at time $t=1$. To construct an $R_{t}$ that generates the probability path $p_{t}$, we seek a rate matrix that satisfies the Kolmogorov equation in Equation 4. Specifically, we learn a model-based rate matrix $R_{t}^{\theta}(y,x)$, parametrised by $\theta$, by minimizing the loss
@@
However, evaluating Equation 5 directly is computationally infeasible due to the intractable summation over $y$, which spans an exponentially large space of possible states, resulting in a complexity of $\mathcal{O}(S^{d})$. To alleviate this issue, we follow Sun et al. (2023c); Campbell et al. (2022); Lou et al. (2024) by assuming independence across dimensions. In particular, we restrict the rate matrix $R_{t}^{\theta}$ such that it assigns non-zero values only to states $y$ that differ from $x$ in at most one dimension. Formally, $R_{t}^{\theta}(y,x)=0$ if $y\notin\mathcal{N}(x)$, where $\mathcal{N}(x):=\{y\in\mathcal{X}|y_{i}\neq x_{i}\ \text{at most one}\ i\}$. To improve clarity in the subsequent sections, we renotate the rate matrix $R_{t}^{\theta}$ for $y\in\mathcal{N}(x)$ with $y_{i}\neq x_{i}$ as
@@
which yields a simplified and more tractable form of the loss in Equation 5
@@
This approximation reduces the computational complexity from $\mathcal{O}(S^{d})$ to $\mathcal{O}(S\times d)$. Nonetheless, two main challenges persist. First, the time derivative $\partial_{t}\log p_{t}(x)$ remains intractable due to the dependence on the partition function, as it expands to $\partial_{t}\log\tilde{p}_{t}(x)-\partial_{t}\log Z_{t}$ with $Z_{t}=\sum_{x}\tilde{p}_{t}(x)$ being intractable. Second, evaluating Equation 7 requires evaluating the neural network $|\mathcal{N}(x)|$ times, which is computationally expensive for each $x$. In the following, we propose several techniques to address these computational bottlenecks.
@@ 3.1 Estimating the Time Derivative of the Log-Partition Function
Figure 1: Comparison of std. dev. and training loss for different estimators of $\partial_{t}\log Z_{t}$. Lower variance estimator exhibits lower training loss, indicating a better learned rate matrix satisfying the Kolmogorov equation in Equation 5.
Because the objective (8) is convex in $c_{t}$, the minimizer is given by $\partial_{t}\log Z_{t}={\mathbb{E}}_{p_{t}}\xi_{t}(x;R_{t})$. Moreover, in practice, the expectation over $p_{t}$ can be safely replaced by an expectation over any distribution $q_{t}$ with the same support, still yielding a valid estimate of the time derivative (see Section A.1 for details). Empirically, we observe that using Equation 8 results in a significantly lower-variance estimator compared to the direct Monte Carlo approach $\partial_{t}\log Z_{t}={\mathbb{E}}_{p_{t}(x)}[\partial_{t}\log\tilde{p}_{t}(x)]$. This reduction in variance can lead to improved optimisation performance. To assess this, we conducted experiments on the Ising model (Mézard et al., 1987), minimising the loss in Equation 7 using two different estimators for $\partial_{t}\log Z_{t}$. The standard deviations of both estimators, as well as their corresponding loss values during training, are plotted over training steps in Figure 1. The results demonstrate that the estimator based on ${\mathbb{E}}_{p_{t}}\xi_{t}(x;R_{t})$ consistently achieves lower loss values, underscoring the benefits of reduced variance in estimating $\partial_{t}\log Z_{t}$ for improved training dynamics. In Section A.2, we provide a perspective of control variate (Geffner & Domke, 2018) to further explain this observation. This insight enables a coordinate descent approach to learning the rate matrix. Specifically: i) $\theta\leftarrow\operatorname*{argmin}_{\theta}\int_{0}^{1}{\mathbb{E}}_{q_{t}(x)}(\xi_{t}(x;R_{t}^{\theta})-c_{t})^{2}\mathop{}\!\mathrm{d}t$; and ii) $c_{t}\leftarrow\operatorname*{argmin}_{c_{t}}{\mathbb{E}}_{q_{t}}(\xi_{t}(x;R_{t})-c_{t})^{2}$. Alternatively, the time derivative can be parameterised directly via a neural network $c_{t}^{\phi}$, allowing joint optimisation of $\theta$ and $\phi$ through the objective $\operatorname*{argmin}_{\theta,\phi}=\int_{0}^{1}{\mathbb{E}}_{q_{t}(x)}(\xi_{t}(x;R_{t}^{\theta})-c_{t}^{\phi})^{2}\mathop{}\!\mathrm{d}t$. This formulation recovers the physics-informed neural network (PINN) loss proposed in Holderrieth et al. (2025). A detailed discussion of the connection to the PINN loss is provided in Section A.3.
@@ 3.2 Efficient Training with Locally Equivariant Networks
As previously noted, computing the $\delta$ function in Equation 7 requires evaluating the neural network $|\mathcal{N}|$ times, which is computationally prohibitive. Inspired by Holderrieth et al. (2025), we proposed to mitigate this issue by utilising locally equivariant networks, an architectural innovation that significantly reduces the computational complexity with the potential to preserve the capacity of network expressiveness. A central insight enabling this reduction is that any rate matrix can be equivalently expressed as a one-way rate matrix [^1] A rate matrix $R$ is said to be one-way if $R(y,x)>0$ implies $R(x,y)=0$. That is, if a transition from $x$ to $y$ is permitted, the reverse transition must be impossible., while still inducing the same probabilistic path. This is formalised in the following proposition:
@@
For a rate matrix $R_{t}$ that generates the probabilistic path $p_{t}$, there exists a one-way rate matrix $Q_{t}(y,x)=\left[R_{t}(y,x)-R_{t}(x,y)\frac{p_{t}(y)}{p_{t}(x)}\right]_{+}$ if $y\neq x$ and $Q_{t}(x,x)=\sum_{y\neq x}Q_{t}(y,x)$, that generates the same probabilistic path $p_{t}$, where $[z]_{+}=\mathrm{max}(z,0)$ denotes the ReLU operation.
This result was originally introduced by Zhang et al. (2023b), and we include a proof in Section B.1 for completeness. Building on Proposition 1, we can parameterise $R_{t}^{\theta}$ directly as a one-way rate matrix. To achieve this, we use a locally equivariant neural network as described by Holderrieth et al. (2025). Specifically, a neural network $G$ is locally equivariant if and only if:
@@
where $\text{Swap}(x,i,\tau)=(x_{1},\dots,x_{i-1},\tau,x_{i+1},\dots,x_{d})$ and $\tau\in\{1,\dots,S\}\triangleq\mathcal{S}$. Based on this, the one-way rate matrix can be defined as $R_{t}^{\theta}(\tau,i|x)\triangleq[G_{t}^{\theta}(\tau,i|x)]_{+}$. Substituting this parametrization into Equation 7, we obtain the simplified expression:
@@
This formulation reduces the computational cost from $\mathcal{O}(|\mathcal{N}|)$ to $\mathcal{O}(1)$, enabling far more efficient training. We term the proposed method as discrete neural flow sampler (DNFS) and summarise the training and sampling details in Appendix C. Nonetheless, the gain in efficiency introduces challenges in constructing a locally equivariant network (leNet) that is both expressive and flexible.
@@ 3.3 Instantiation of leNets: Locally Equivariant Transformer
To construct a locally equivariant network, we first introduce hollow network (Chen & Duvenaud, 2019). Formally, let $x=(x_{1},\dots,x_{i}=\tau,\dots,x_{d})$ denote the input with its $i$-th token set to $\tau$. A function $H:\mathcal{X}\mapsto\mathbb{R}^{d\times h}$ is termed a hollow network if it satisfies $H(x_{i\leftarrow\tau})_{i,:}=H(x_{i\leftarrow\tau^{\prime}})_{i,:},\forall\tau,\tau^{\prime}\in\mathcal{S}$, where $M_{i,:}$ denotes the $i$-th row of the matrix $M$. Intuitively, it implies that the output at position $i$ is invariant to the value of the $i$-th input token. Hollow networks provide a foundational building block for constructing locally equivariant networks, as formalised in the following proposition.
@@
This can be verified via $G(\tau,i|x)\!=\!-(\omega_{x_{i}}\!-\!\omega_{\tau})^{T}H(\text{Swap}(x,i,\tau))_{i,:}\!=\!-G(x_{i},i|\text{Swap}(x,i,\tau))$. Although Proposition 2 offers a concrete approach to constructing locally equivariant networks, significant challenges persist. In contrast to globally equivariant architectures (Cohen & Welling, 2016; Fuchs et al., 2020), where the composition of equivariant layers inherently preserves equivariance, locally equivariant networks are more delicate to design. In particular, stacking locally equivariant layers does not, in general, preserve local equivariance. While Holderrieth et al. (2025) propose leveraging multi-layer perceptions (MLPs), attention mechanisms, and convolutional layers (see Section B.2 for details) to construct locally equivariant networks, these architectures may still fall short in terms of representational capacity and flexibility.
@@
Locally Equivariant Transformer. As shown in Proposition 2, a crucial component to construct a locally equivariant network is the hollow network. Specifically, the key design constraint is that the network’s output at dimension $i$ must be independent of the corresponding input value $x_{i}$. Otherwise, any dependence would result in information leakage and violate local equivariance. However, the $i$-th output may depend freely on all other coordinates of the input token $x$ except for the $i$-th entry. This insight motivates the use of hollow transformers (Sun et al., 2023c) as a foundation for constructing locally equivariant networks. Specifically, it employs two autoregressive Transformers (Vaswani et al., 2017; Radford et al., 2018) per layer; one processing inputs from left to right, and the other from right to left. In the readout layer, the representations from two directions are fused via attention to produce the output. This design ensures that each output dimension remains independent of its corresponding input coordinate, while still leveraging the expressiveness of multi-layer Transformers. Thereby, the final output $G_{t}^{\theta}(\cdot,i|x)$ can be obtained by taking the inner product between the hollow attention output and the token embeddings produced by the projection layer. We term the proposed architecture as locally equivariant transformer (leTF), and defer the implementation details to Section B.3.
Figure 3: Comparison of log RMSE ()($\downarrow$)( ↓ ) and ESS ()($\uparrow$)( ↑ ) for different locally equivariant networks. More expressive networks achieve better performance.
Comparison of Different leNets. In Figure 3, we compare leTF with other locally equivariant networks for training a discrete neural flow sampler on the Ising model. The results show that leTF achieves lower estimation errors and higher effective sample sizes (see Section E.1 for experimental details). leAttn and leMLP, which each consist of only a single locally equivariant layer, perform significantly worse, highlighting the importance of network expressiveness in achieving effective local equivariance. Although leConv performs comparably to leTF, its convolutional design is inherently less flexible. It is restricted to grid-structured data, such as images or the Ising model, and does not readily generalise to other data types like text or graphs. Additionally, as shown in Figure 15, leTF achieves lower training loss compared to leConv, further confirming its advantage in expressiveness.

### v2

@@ 1 Introduction
Building on these advances, the goal of our paper is to develop a sampling method for discrete distributions that is both efficient and scalable. To this end, we introduce Discrete Neural Flow Samplers (DNFS), a novel framework that learns the rate matrix of a CTMC whose dynamics satisfy the Kolmogorov forward equation (Oksendal, 2013). In contrast to discrete flow models (Campbell et al., 2024; Gat et al., 2024), which benefit from access to training data to fit the generative process, DNFS operates in settings where no data samples are available. This data-free setting makes direct optimisation of the Kolmogorov objective particularly challenging and necessitates new methodological advances to ensure stable and effective training. Specifically, the first difficulty lies in the dependence of the objective on the intractable partition function. We mitigate this by using control variates (Geffner & Domke, 2018) to reduce the variance of its Monte Carlo estimate, which enables efficient optimisation via coordinate descent. More critically, standard neural network parameterisations of the rate matrix render the objective computationally prohibitive. To make learning tractable, a locally equivariant Transformer architecture is introduced to enhance computational efficiency significantly while retaining strong model expressiveness. Empirically, DNFS proves to be an effective sampler for discrete unnormalised distributions. We further demonstrate its versatility in diverse applications, including training discrete energy-based models and solving combinatorial optimization problems.
@@ 2 Preliminaries
In the next section, we describe how to leverage Equation˜4 to learn a model-based rate matrix for sampling from a given target distribution $\pi$, followed by the discussion of applications to discrete energy-based modelling and combinatorial optimisation.
@@ 3 Discrete Neural Flow Samplers
The rate matrix $R_{t}$ that transports the initial distribution to the target is generally not unique. However, we can select a particular path by adopting an annealing interpolation between the prior $\eta$ and the target $\rho$, defined as $p_{t}\propto\rho^{t}\eta^{1-t}\triangleq\tilde{p}_{t}$ (Gelman & Meng, 1998; Neal, 2001). This annealing path coincides with the target distribution $\pi\propto\rho$ at time $t=1$. To construct an $R_{t}$ that generates the probability path $p_{t}$, we seek a rate matrix that satisfies the Kolmogorov equation in Equation˜4. Specifically, we learn a model-based rate matrix $R_{t}^{\theta}(y,x)$, parametrised by $\theta$, by minimizing the loss
@@
However, evaluating Equation˜5 directly is computationally infeasible due to the intractable summation over $y$, which spans an exponentially large space of possible states, resulting in a complexity of $\mathcal{O}(S^{d})$. To alleviate this issue, we follow Sun et al. (2023c); Campbell et al. (2022); Lou et al. (2024) by assuming independence across dimensions. In particular, we restrict the rate matrix $R_{t}^{\theta}$ such that it assigns non-zero values only to states $y$ that differ from $x$ in at most one dimension. Formally, $R_{t}^{\theta}(y,x)=0$ if $y\notin\mathcal{N}(x)$, where $\mathcal{N}(x):=\{y\in\mathcal{X}|y_{i}\neq x_{i}\ \text{at most one}\ i\}$. To improve clarity in the subsequent sections, we renotate the rate matrix $R_{t}^{\theta}$ for $y\in\mathcal{N}(x)$ with $y_{i}\neq x_{i}$ as
@@
which yields a simplified and more tractable form of the loss in Equation˜5
@@
This approximation reduces the computational complexity from $\mathcal{O}(S^{d})$ to $\mathcal{O}(S\times d)$. Nonetheless, two main challenges persist. First, the time derivative $\partial_{t}\log p_{t}(x)$ remains intractable due to the dependence on the partition function, as it expands to $\partial_{t}\log\tilde{p}_{t}(x)-\partial_{t}\log Z_{t}$ with $Z_{t}=\sum_{x}\tilde{p}_{t}(x)$ being intractable. Second, evaluating Equation˜7 requires evaluating the neural network $|\mathcal{N}(x)|$ times, which is computationally expensive for each $x$. In the following, we propose several techniques to address these computational bottlenecks.
@@ 3.1 Estimating the Time Derivative of the Log-Partition Function
Figure 1: Comparison of std. dev. and training loss for different estimators of $\partial_{t}\log Z_{t}$. Lower variance estimator exhibits lower training loss, indicating a better learned rate matrix satisfying the Kolmogorov equation in Equation˜5.
Because the objective (8) is convex in $c_{t}$, the minimizer is given by $\partial_{t}\log Z_{t}={\mathbb{E}}_{p_{t}}\xi_{t}(x;R_{t})$. Moreover, in practice, the expectation over $p_{t}$ can be safely replaced by an expectation over any distribution $q_{t}$ with the same support, still yielding a valid estimate of the time derivative (see Section˜A.1 for details). Empirically, we observe that using Equation˜8 results in a significantly lower-variance estimator compared to the direct Monte Carlo approach $\partial_{t}\log Z_{t}={\mathbb{E}}_{p_{t}(x)}[\partial_{t}\log\tilde{p}_{t}(x)]$. This reduction in variance can lead to improved optimisation performance. To assess this, we conducted experiments on the Ising model (Mézard et al., 1987), minimising the loss in Equation˜7 using two different estimators for $\partial_{t}\log Z_{t}$. The standard deviations of both estimators, as well as their corresponding loss values during training, are plotted over training steps in Figure˜1. The results demonstrate that the estimator based on ${\mathbb{E}}_{p_{t}}\xi_{t}(x;R_{t})$ consistently achieves lower loss values, underscoring the benefits of reduced variance in estimating $\partial_{t}\log Z_{t}$ for improved training dynamics. In Section˜A.2, we provide a perspective of control variate (Geffner & Domke, 2018) to further explain this observation. This insight enables a coordinate descent approach to learning the rate matrix. Specifically: i) $\theta\leftarrow\operatorname*{argmin}_{\theta}\int_{0}^{1}{\mathbb{E}}_{q_{t}(x)}(\xi_{t}(x;R_{t}^{\theta})-c_{t})^{2}\mathop{}\!\mathrm{d}t$; and ii) $c_{t}\leftarrow\operatorname*{argmin}_{c_{t}}{\mathbb{E}}_{q_{t}}(\xi_{t}(x;R_{t})-c_{t})^{2}$. Alternatively, the time derivative can be parameterised directly via a neural network $c_{t}^{\phi}$, allowing joint optimisation of $\theta$ and $\phi$ through the objective $\operatorname*{argmin}_{\theta,\phi}=\int_{0}^{1}{\mathbb{E}}_{q_{t}(x)}(\xi_{t}(x;R_{t}^{\theta})-c_{t}^{\phi})^{2}\mathop{}\!\mathrm{d}t$. This formulation recovers the physics-informed neural network (PINN) loss proposed in Holderrieth et al. (2025). A detailed discussion of the connection to the PINN loss is provided in Section˜A.3.
@@ 3.2 Efficient Training with Locally Equivariant Networks
As previously noted, computing the $\delta$ function in Equation˜7 requires evaluating the neural network $|\mathcal{N}|$ times, which is computationally prohibitive. Inspired by Holderrieth et al. (2025), we proposed to mitigate this issue by utilising locally equivariant networks, an architectural innovation that significantly reduces the computational complexity with the potential to preserve the capacity of network expressiveness. A central insight enabling this reduction is that any rate matrix can be equivalently expressed as a one-way rate matrix [^1] A rate matrix $R$ is said to be one-way if $R(y,x)>0$ implies $R(x,y)=0$. That is, if a transition from $x$ to $y$ is permitted, the reverse transition must be impossible., while still inducing the same probabilistic path. This is formalised in the following proposition:
@@
For a rate matrix $R_{t}$ that generates the probabilistic path $p_{t}$, there exists a one-way rate matrix $Q_{t}(y,x)=\left[R_{t}(y,x)-R_{t}(x,y)\frac{p_{t}(y)}{p_{t}(x)}\right]_{+}$ if $y\neq x$ and $Q_{t}(x,x)=\sum_{y\neq x}Q_{t}(y,x)$, that generates the same probabilistic path $p_{t}$, where $z]_{+}=\mathrm{max}(z,0)$ denotes the ReLU operation.
This result was originally introduced by Zhang et al. (2023b), and we include a proof in Section˜B.1 for completeness. Building on Proposition˜1, we can parameterise $R_{t}^{\theta}$ directly as a one-way rate matrix. To achieve this, we use a locally equivariant neural network as described by Holderrieth et al. (2025). Specifically, a neural network $G$ is locally equivariant if and only if:
@@
where $\text{Swap}(x,i,\tau)=(x_{1},\dots,x_{i-1},\tau,x_{i+1},\dots,x_{d})$ and $\tau\in\{1,\dots,S\}\triangleq\mathcal{S}$. Based on this, the one-way rate matrix can be defined as $R_{t}^{\theta}(\tau,i|x)\triangleq[G_{t}^{\theta}(\tau,i|x)]_{+}$. Substituting this parametrization into Equation˜7, we obtain the simplified expression:
@@
This formulation reduces the computational cost from $\mathcal{O}(|\mathcal{N}|)$ to $\mathcal{O}(1)$, enabling far more efficient training. We term the proposed method as discrete neural flow sampler (DNFS) and summarise the training and sampling details in Appendix˜C. Nonetheless, the gain in efficiency introduces challenges in constructing a locally equivariant network (leNet) that is both expressive and flexible.
@@ 3.3 Instantiation of leNets: Locally Equivariant Transformer
To construct a locally equivariant network, we first introduce hollow network (Chen & Duvenaud, 2019). Formally, let $x_{i\leftarrow\tau}\!=\!(x_{1},\dots,x_{i}\!=\!\tau,\dots,x_{d})$ denote the input with its $i$-th token set to $\tau$. A function $H:\mathcal{X}\mapsto\mathbb{R}^{d\times h}$ is termed a hollow network if it satisfies $H(x_{i\leftarrow\tau})_{i,:}=H(x_{i\leftarrow\tau^{\prime}})_{i,:},\forall\tau,\tau^{\prime}\in\mathcal{S}$, where $M_{i,:}$ denotes the $i$-th row of the matrix $M$. Intuitively, it implies that the output at position $i$ is invariant to the value of the $i$-th input token. Hollow networks provide a foundational building block for constructing locally equivariant networks, as formalised in the following proposition.
@@
This can be verified via $G(\tau,i|x)\!=\!-(\omega_{x_{i}}\!-\!\omega_{\tau})^{T}H(\text{Swap}(x,i,\tau))_{i,:}\!=\!-G(x_{i},i|\text{Swap}(x,i,\tau))$. Although Proposition˜2 offers a concrete approach to constructing locally equivariant networks, significant challenges persist. In contrast to globally equivariant architectures (Cohen & Welling, 2016; Fuchs et al., 2020), where the composition of equivariant layers inherently preserves equivariance, locally equivariant networks are more delicate to design. In particular, stacking locally equivariant layers does not, in general, preserve local equivariance. While Holderrieth et al. (2025) propose leveraging multi-layer perceptions (MLPs), attention mechanisms, and convolutional layers (see Section˜B.2 for details) to construct locally equivariant networks, these architectures may still fall short in terms of representational capacity and flexibility.
@@
Locally Equivariant Transformer. As shown in Proposition˜2, a crucial component to construct a locally equivariant network is the hollow network. Specifically, the key design constraint is that the network’s output at dimension $i$ must be independent of the corresponding input value $x_{i}$. Otherwise, any dependence would result in information leakage and violate local equivariance. However, the $i$-th output may depend freely on all other coordinates of the input token $x$ except for the $i$-th entry. This insight motivates the use of hollow transformers (Sun et al., 2023c) as a foundation for constructing locally equivariant networks. Specifically, it employs two autoregressive Transformers (Vaswani et al., 2017; Radford et al., 2018) per layer; one processing inputs from left to right, and the other from right to left. In the readout layer, the representations from two directions are fused via attention to produce the output. This design ensures that each output dimension remains independent of its corresponding input coordinate, while still leveraging the expressiveness of multi-layer Transformers. Thereby, the final output $G_{t}^{\theta}(\cdot,i|x)$ can be obtained by taking the inner product between the hollow attention output and the token embeddings produced by the projection layer. We term the proposed architecture as locally equivariant transformer (leTF), and defer the implementation details to Section˜B.3.
Figure 3: Comparison of log RMSE $(\downarrow)$ and ESS $(\uparrow)$ for different locally equivariant networks. More expressive networks achieve better performance.
Comparison of Different leNets. In Figure˜3, we compare leTF with other locally equivariant networks for training a discrete neural flow sampler on the Ising model. The results show that leTF achieves lower estimation errors and higher effective sample sizes (see Section˜E.1 for experimental details). leAttn and leMLP, which each consist of only a single locally equivariant layer, perform significantly worse, highlighting the importance of network expressiveness in achieving effective local equivariance. Although leConv performs comparably to leTF, its convolutional design is inherently less flexible. It is restricted to grid-structured data, such as images or the Ising model, and does not readily generalise to other data types like text or graphs. Additionally, as shown in Figure˜15, leTF achieves lower training loss compared to leConv, further confirming its advantage in expressiveness.